OpenStudy (anonymous):

Describe how to transform the graph of f into the graph of g. f(x) = x^3 and g(x) = -(-x^3)

OpenStudy (anonymous):

Reflect the graph of f across the x-axis.

OpenStudy (ghazi):

wait -(-x^3)= x^3= f(x)

OpenStudy (anonymous):

Describe how to transform the graph of f into the graph of g. f(x) = x^3 and g(x) = -(-x^3)

OpenStudy (ghazi):

first of all take the mirror image of graph of f(x) you will get -f(x)

OpenStudy (anonymous):

ok so reflect over y axis

OpenStudy (ghazi):

yes, and twice

OpenStudy (anonymous):

Reflect the graph of f across the x-axis and then reflect across the y-axis or just over Y?

OpenStudy (ghazi):

just y

OpenStudy (ghazi):